An RLC circuit is connected to 60hz, 240volts a.c supply. If R=50ohms, C=100uf and M=0.12h. Calculate the power used in circuit


1
Expert's answer
2021-02-22T10:26:18-0500
"P=\frac{U^2}{Z}\\P=\frac{U^2}{\sqrt{R^2+(2\pi f L-\frac{1}{2\pi f C})^2}}
\\P=\frac{240^2}{\sqrt{50^2+(2\pi (60)(0.12)-\frac{1}{2\pi (60)(10^{-4})})^2}}=1079\ W"